OpenStudy (anonymous):

if A=22.5 than (1+cosA)(1+cos3A)(1+cos5A)(1+cos7A)=?

OpenStudy (anonymous):

no one replyyyyy?????

OpenStudy (anonymous):

Are you alllowed to use a calculator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

bandar kya jane jao madari ke paas

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

Where you are stuck

OpenStudy (anonymous):

Reply

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

have you joined school-classXI

OpenStudy (anonymous):

i am preparing for IIT , i am in 11th

OpenStudy (anonymous):

Reply

OpenStudy (anonymous):

??? go for trigo SLloney

OpenStudy (anonymous):

Lol, I can do this problem orally , I know how to solve these problems

OpenStudy (anonymous):

then write

OpenStudy (anonymous):

U have A =22.5 So u can calculate 3A ,5A , 7A ?

OpenStudy (anonymous):

every one a small kid of class 5 know think How will u crack IIT I need answer not ideas of nursery

OpenStudy (anonymous):

Okay so if you are allowed to use a calculator It equals, (1+cos22.5)(1+cos67.5)(1+cos112.5)(1+cos157.5) Answer comes to be 0.125

OpenStudy (anonymous):

(1+cos22.5)(1+cos67.5)(1+cos112.5)(1+cos157.5) value? without using calculator its class 11 Q

OpenStudy (anonymous):

challange for u like IIT

OpenStudy (anonymous):

cos(45)= 1/√2 cos(θ/2) = +/-sqrt [(1+cos(θ))/2] cos(22.5) = +/- sqrt {1+cos(45)) /2] = +/- sqrt [ (1+1/√2) /2] = +/- sqrt [ (√2+1) / 2√2)] Like that you can use the half-angle identity in trigonometry Or i can give you another method (1+cos22.5) (1+cos67.5)(1+cos 112.5)(1+cos 157.5)? =(1+cos22.5) {1+cos(90−22.5)}{1+cos(90+22.5)} {1+cos(180−22.5)} =(1+cos22.5) {1+sin22.5)}{1−sin22.5}{1−cos22.5} =(1+cos22.5){1−cos22.5} {1+sin22.5}{1−sin22.5} ={1−cos²22.5} {1−sin²22.5} ={sin²22.5 cos²22.5} ={sin22.5 cos22.5}² ={(sin45)/2}² ={(1/√2)/2}²=1/8

OpenStudy (anonymous):

best answer

OpenStudy (anonymous):

Thank you! you are welcome

OpenStudy (anonymous):

I thought you have to do with calculator , which is very dumb that's why i was trolling nvm

OpenStudy (anonymous):

Which class are you in

OpenStudy (anonymous):

You have blocked me

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Its time for u to prepare only for IIT not for chatting thats why I block you bye over

OpenStudy (anonymous):

LOL, okay Which class are you in

OpenStudy (anonymous):

Bye its close all the best but behind me look on 2016 result my name in the TOPPER & YOU ARE FOLLWER sorry &bye

OpenStudy (anonymous):

What ?????